# CodeMirage: Hallucinations in Code Generated by Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have shown promising potentials in program generation and no-code automation. However, LLMs are prone to generate hallucinations, i.e., they generate text which sounds plausible but is incorrect. Although there has been a recent surge in research on LLM hallucinations for text generation, similar hallucination phenomenon can happen in code generation. Sometimes the generated code can have syntactical or logical errors as well as more advanced issues like security vulnerabilities, memory leaks, etc. Given the wide adaptation of LLMs to enhance efficiency in code generation and development in general, it becomes imperative to investigate hallucinations in code generation. To the best of our knowledge, this is the first attempt at studying hallucinations in the code generated by LLMs. We start by introducing the code hallucination definition and a comprehensive taxonomy of code hallucination types. We propose the first benchmark CodeMirage dataset [^1] The dataset will be released upon acceptance. for code halllucinations. The benchmark contains 1,13711371,1371 , 137 GPT-3.5 generated hallucinated code snippets for Python programming problems from two base datasets – HumanEval and MBPP. We then propose the methodology for code hallucination detection and experiment with open source LLMs such as CodeLLaMA as well as OpenAI’s GPT-3.5 and GPT-4 models using one-shot prompt. We find that GPT-4 performs the best on HumanEval dataset and gives comparable results to the fine-tuned CodeBERT baseline on MBPP dataset. Towards the end, we discuss various mitigation strategies for code hallucinations and conclude our work.
@@ 2.2 Hallucinations in LLMs
Although LLMs have shown remarkable performance in natural language generation, they still inevitably encounter several issues, hallucination being one of the top Wu et al. (2024); Ghafouri et al. (2023); Huang et al. (2023). Ji et al. (2023) defined hallucination as the generated content that is nonsensical or unfaithful to the provided source content. Previous works Ji et al. (2023); Maynez et al. (2020); Huang et al. (2021) categorized hallucination into two main categories – intrinsic and extrinsic. Intrinsic hallucination happens when the generated output contradicts the source content, whereas extrinsic hallucination happens when the generated output cannot be verified from the source content (i.e., output that can neither be supported nor contradicted by the source). Within the context of LLMs, Zhang et al. (2023b) defined hallucination into 3 categories – input-conflicting (LLM-generated content deviates from the source input provided by users), context-conflicting (LLM-generated content conflicts with previously generated information by itself), and text-conflicting (LLM-generated content is not faithful to the established world knowledge) hallucinations.

### v2

@@ Abstract
Large Language Models (LLMs) have shown promising potentials in program generation and no-code automation. However, LLMs are prone to generate hallucinations, i.e., they generate text which sounds plausible but is incorrect. Although there has been a recent surge in research on LLM hallucinations for text generation, similar hallucination phenomenon can happen in code generation. Sometimes the generated code can have syntactical or logical errors as well as more advanced issues like security vulnerabilities, memory leaks, etc. Given the wide adaptation of LLMs to enhance efficiency in code generation and development in general, it becomes imperative to investigate hallucinations in code generation. To the best of our knowledge, this is the first attempt at studying hallucinations in the code generated by LLMs. We start by introducing the code hallucination definition and a comprehensive taxonomy of code hallucination types. We propose the first benchmark CodeMirage dataset [^1] The dataset will be released upon acceptance. for code hallucinations. The benchmark contains 1,13711371,1371 , 137 GPT-3.5 generated hallucinated code snippets for Python programming problems from two base datasets – HumanEval and MBPP. We then propose the methodology for code hallucination detection and experiment with open source LLMs such as CodeLLaMA as well as OpenAI’s GPT-3.5 and GPT-4 models using one-shot prompt. We find that GPT-4 performs the best on HumanEval dataset and gives comparable results to the fine-tuned CodeBERT baseline on MBPP dataset. Towards the end, we discuss various mitigation strategies for code hallucinations and conclude our work.
@@ 2.2 Hallucinations in LLMs
Although LLMs have shown remarkable performance in natural language generation, they still inevitably encounter several issues, hallucination being one of the top Wu et al. (2024); Ghafouri et al. (2023); Huang et al. (2023). Ji et al. (2023) defined hallucination as the generated content that is nonsensical or unfaithful to the provided source content. Previous works Ji et al. (2023); Maynez et al. (2020); Huang et al. (2021) categorized hallucination into two main categories – intrinsic and extrinsic. Intrinsic hallucination happens when the generated output contradicts the source content, whereas extrinsic hallucination happens when the generated output cannot be verified from the source content (i.e., output that can neither be supported nor contradicted by the source). Within the context of LLMs, Zhang et al. (2023b) defined hallucination into 3 categories – input-conflicting (LLM-generated content deviates from the source input provided by users), context-conflicting (LLM-generated content conflicts with previously generated information by itself), and fact-conflicting (LLM-generated content is not faithful to the established world knowledge) hallucinations.
@@ 6 Conclusions and Future Work
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.